# Evaluating, Synthesizing, and Enhancing for Customer Support Conversation

### v1

@@ 1 Introduction
Customer support aims to help users resolve product- or service-related issues through effective and context-aware communication. While large-scale dialogue systems have received growing attention in recent years (Budzianowski et al. 2018; Rashkin et al. 2019; Peskov et al. 2019; Liu et al. 2021), customer support conversations remain underexplored in the NLP community, [^1] This paper uses dialogue and conversation interchangeably. largely due to the scarcity of publicly available benchmarks and the sensitive, domain-specific nature of support interactions. Meanwhile, large language models (LLMs) like GPT-3 (Brown et al. 2020) have shown impressive capabilities in open-domain dialogue generation (Zhang et al. 2020b; Bae et al. 2022; Thoppilan et al. 2022), but their ability to generate realistic and effective customer support conversations remains underexamined.
@@
Customer support guidelines from COPC, the internationally recognized standard for customer experience management, [^2] https://www.copc.com/copc-standards/ emphasize that high-quality support often depends on structured communication strategies, similar to those used in emotional support settings (Liu et al. 2021). As shown in Figure 1, [^3] The example is translated into English for better readability. the service supporter begins with the Greeting strategy to initiate the conversation and explore the customer’s issue. Upon recognizing the customer’s negative emotion, the supporter adopts the Emotional Management strategy, expressing empathy and understanding to alleviate the customer’s distress. After understanding the problem, the supporter employs the Information Delivery strategy to provide clear and actionable guidance, followed by the Feedback Request strategy to check for further concerns. Finally, the conversation ends with the Appreciation and Closure strategy, ensuring a respectful and reassuring followed by Feedback Request to check for further concerns.
@@
Despite the importance of effective customer support, research on real-time customer service conversations remains limited, mainly due to a lack of task-specific design and high-quality annotated data. Many studies (Xu et al. 2017; Oraby et al. 2017; Cui et al. 2017; Mesquita, Martins, and Almeida 2022) rely on asynchronous, exchanges (e.g., Twitter), where interactions span minutes to days. These settings differ significantly from the immediate, uninterrupted nature of real-time support. Moreover, effective support requires not only resolving issues but also showing empathy and emotional support. Yet, most task-oriented dialogue datasets (Wen et al. 2017; Budzianowski et al. 2018; Peskov et al. 2019; Gung et al. 2023) lack the intentional use of supportive strategies like emotional management or empathetic closure, which are vital for high-quality customer service.
@@
To address this gap, we introduce the task of customer support conversation (CSC), to facilitate the training of customer service supporters. The goal is to help them respond with appropriate strategies that combine accurate solutions with empathetic communication. Based on COPC standards and inspired by the emotional support framework (Liu et al. 2021), we develop a CSC-specific framework with five dialogue stages and twelve support strategies. Using it, we build CSConv, a high-quality dataset adapted from real service interactions and refined for structured strategy use. To address the lack of high-quality training data, we also develop a role-playing approach that creates RoleCS, a synthetic dataset of strategy-rich conversations by assigning LLMs distinct roles aligned with the CSC framework. Fine-tuning on RoleCS significantly boosts LLMs’ ability to generate strategy-aligned and effective responses on CSConv.
@@ 2 Related Work
The goal of CSConv, which is to generate conversations based on real spoken customer service interactions, differs significantly from previous task-oriented dialogue datasets. Many earlier synthetic datasets have been created using the Wizard of Oz (WOZ) framework (Kelley 1984), where one person acts as the system and another as the user. Wen et al. (2017) introduce a crowdsourced version of the WOZ framework to collect domain-specific dialogue data more efficiently. Other task-oriented dialogue datasets include Frames (El Asri et al. 2017), MultiWOZ (Budzianowski et al. 2018), MultiDoGO (Peskov et al. 2019), EMPATHETICDIALOGUES (Rashkin et al. 2019), Taskmaster-1 (Byrne et al. 2019), ESConv (Liu et al. 2021), and NATCS (Gung et al. 2023), each of which explores different domains and annotation schemes for developing conversational agents.
@@ 2.2 Role-Playing in Conversation Generation
Recent advances in LLMs have enabled the use of role-playing agents for conversation generation. For example, Bae et al. (2022) generate dialogues aligned with specific roles, while Yang et al. (2024b) simulate characters and settings to produce coherent interactions. Wu et al. (2024) create interactive narrative dramas, and Ye et al. (2025) focus on emotional support conversations. Building on the Ye et al. (2025), we design role-playing agents for customer support with clearer role definitions and responsibilities to more closely reflect real-world service interactions.
@@ 3.1 CSC Framework
Building on the three core stages of supportive communication from Hill (2019) and the COPC practical guidelines, we work with domain experts to refine and extend this structure for customer support. The resulting CSC framework defines five stages: Connecting (greeting and rapport building), Identifying (understanding the customer’s issue and emotional state), Exploring (discussing and evaluating potential solutions), Resolving (delivering and confirming resolution), and Maintaining (closing the interaction while preserving the customer relationship). Figure 2 illustrates the flow across these stages.
@@ 3.2 Dataset Construction
Appendix B provides details on filtering rules, prompt templates, and annotation guidelines. For rewriting, we use DeepSeek-R1 (Guo et al. 2025) instead of GPT-4o (OpenAI 2024), as the latter tends to produce shorter, less emotionally rich dialogues. See Appendix B.5 for a comparison.
@@ 3.4 CSC Task Definition
Following the task of emotional support conversation (Liu et al. 2021; Ye et al. 2025), we define the CSC task as generating the supporter’s response. Specifically, at turn $k$, where $P_{k}=S$, the model receives the conversation history $X_{k}=\{\left(P_{i},T_{i},U_{i}\right)\}_{i=1}^{k-1}$ as input. The CSC task is then divided into two subtasks: [^5] We further investigate how strategy prediction contributes to improving response generation in Section 6.3.
@@ 4 Synthetic Conversation Generation with Role-Playing Agents
Fine-tuning LLMs has shown effective for improving performance in task-specific dialogue applications. However, building high-quality, multi-turn customer support datasets remains challenging due to the need for domain expertise, contextual coherence, and dialogue diversity. This limits both the scalability and scenario coverage. While recent studies use LLMs for dataset augmentation via rewriting or imitation, they often yield limited variation in dialogue flow and support strategy. To address this, inspired by Ye et al. (2025), we adopt a multi-role role-playing framework to generate synthetic customer support dialogues that are diverse, coherent, and representative of real-world scenarios.
@@ 4.1 Role Playing Conversation Generation
As shown in Figure 4, our role-playing framework involves five roles: Planner, Supporter Assistant, Supporter, Customer Assistant, and Customer. Each role serves a distinct purpose. The Planner defines the dialogue scenario and sets the Customer’s communication goal. The main interaction occurs between the Supporter and the Customer, simulating a real customer service exchange. Meanwhile, their respective Assistants, the Supporter Assistant and the Customer Assistant, offer strategic guidance to help them fulfill their roles more effectively. We use deepseek-r1 (Guo et al. 2025) to simulate all roles, with detailed prompts available in Appendix E. We describe each role below.
@@
Diversity in conversations relies heavily on the richness of customer personas (Wang et al. 2025). To support this, we construct a rich character profile pool $O$ that captures a wide spectrum of customer backgrounds, behaviors, and communication styles, thereby enhancing the realism and variability of the simulated interactions.
@@ 5.1 Experimental Settings
We evaluate several widely used LLMs on the CSC task, including GPT-4o (OpenAI 2024), DeepSeek-R1 (Guo et al. 2025), DeepSeek-V3 (Liu et al. 2024), Qwen-2.5-7B/72B-Instruct (Yang et al. 2024a), and LLaMA-3.1-8B/70B-Instruct (Grattafiori et al. 2024). To evaluate the impact of the proposed RoleCS, we fine-tune the Qwen and LLaMA models using 137,406 fine-tuning instances extracted from RoleCS, formatted consistently with our single-prompt in test-time. Fine-tuning and inference configurations are provided in Appendix J.
@@
Following prior work (Liu et al. 2021; Ye et al. 2025), we adopt a diverse set of metrics to assess the response quality. These include BLEU-n (B-n(Papineni et al. 2002) and ROUGE-L (R-L(Lin 2004) for measuring lexical overlap, BERTScore (BS(Zhang et al. 2020a) and BLEURT (BR(Sellam, Das, and Parikh 2020) for semantic similarity. In addition, to evaluate alignment with the intended support strategy, which is a core aspect of the CSC task, we report strategy prediction accuracy (ACC).
@@ 5.2 Experimental Results
  • Evaluation with generated context shows similar performance trends as with reference context, but with lower absolute scores. This drop reflects the difficulty of maintaining consistency and quality in multi-turn conversations when relying on model-generated dialogue history, echoing findings in prior work such as Ye et al. (2025).
@@ 6.4 Evaluation with LLMs and Human as Judges
For human evaluation, we randomly select 100 conversations and have them independently rated by three professional annotators on a 1-5 Likert scale (Joshi et al. 2015) across the same six dimensions. Table 6 shows the average overall scores, which align with trends from the LLM-based evaluations. Table 7 reports Fleiss’ Kappa (Fleiss 1971) scores among annotators and between GPT-4o-Judge and the annotators. The results show strong inter‑rater agreement and confirm that GPT-4o-Judge aligns well with human judgment.
@@ Appendix H Diversity Analysis
We examine the diversity of the CSConv and RoleCS datasets from two perspectives: lexical diversity and semantic diversity. To assess lexical diversity, we adopt Distinct-n (Li et al. 2016) metric. For a fair comparison, we randomly select 1,000 dialogues from each dataset, removing speaker and strategy information, and utilizing Qwen2.5-7B tokenizer for tokenization. Regarding semantic diversity, we calculate cosine similarity between pairs of distinct dialogues using TF-IDF features (Salton, Wong, and Yang 1975). Table 12 and Figure 28 present the results, respectively.
@@ Appendix I Prompt for Predicting Strategy and Generating Response
For fine-tuning, we use the LLaMA-Factory framework (Zheng et al. 2024), applying LoRA (Hu et al. 2022) with a rank of 8 and a scaling factor of 16. All fine-tunings are conducted on 4 NVIDIA A100 80GB GPUs, with a batch size of 4 and gradient accumulation over 2 steps. The learning rate is set to 3e-5, and fine-tuning is run for 3 epochs to mitigate the risk of overfitting.

### v2

@@ 1 Introduction
Customer support aims to help users resolve product- or service-related issues through effective and context-aware communication. While large-scale dialogue systems have received growing attention in recent years (budzianowski-etal-2018-multiwoz; rashkin-etal-2019-towards; peskov-etal-2019-multi; liu-etal-2021-towards), customer support conversations remain underexplored in the NLP community, [^1] This paper uses dialogue and conversation interchangeably. largely due to the scarcity of publicly available benchmarks and the sensitive, domain-specific nature of support interactions. Meanwhile, large language models (LLMs) like GPT-3 (brown-etal-2020-language) have shown impressive capabilities in open-domain dialogue generation (zhang-etal-2020-dialogpt; bae-etal-2022-building; thoppilan-etal-2022-lamda), but their ability to generate realistic and effective customer support conversations remains underexamined.
@@
Customer support guidelines from COPC, the internationally recognized standard for customer experience management, [^2] https://www.copc.com/copc-standards/ emphasize that high-quality support often depends on structured communication strategies, similar to those used in emotional support settings (liu-etal-2021-towards). As shown in Figure 1, [^3] The example is translated into English for better readability. the service supporter begins with the Greeting strategy to initiate the conversation and explore the customer’s issue. Upon recognizing the customer’s negative emotion, the supporter adopts the Emotional Management strategy, expressing empathy and understanding to alleviate the customer’s distress. After understanding the problem, the supporter employs the Information Delivery strategy to provide clear and actionable guidance, followed by the Feedback Request strategy to check for further concerns. Finally, the conversation ends with the Appreciation and Closure strategy, ensuring a respectful and reassuring followed by Feedback Request to check for further concerns.
@@
Despite the importance of effective customer support, research on real-time customer service conversations remains limited, mainly due to a lack of task-specific design and high-quality annotated data. Many studies (xu-etal-2017-a; oraby-etal-2017-how; cui-etal-2017-superagent; mesquita-etal-2022-dense) rely on asynchronous, exchanges (e.g., Twitter), where interactions span minutes to days. These settings differ significantly from the immediate, uninterrupted nature of real-time support. Moreover, effective support requires not only resolving issues but also showing empathy and emotional support. Yet, most task-oriented dialogue datasets (wen-etal-2017-network; budzianowski-etal-2018-multiwoz; peskov-etal-2019-multi; gung-etal-2023-natcs) lack the intentional use of supportive strategies like emotional management or empathetic closure, which are vital for high-quality customer service.
@@
To address this gap, we introduce the task of customer support conversation (CSC), to facilitate the training of customer service supporters. The goal is to help them respond with appropriate strategies that combine accurate solutions with empathetic communication. Based on COPC standards and inspired by the emotional support framework (liu-etal-2021-towards), we develop a CSC-specific framework with five dialogue stages and twelve support strategies. Using it, we build CSConv, a high-quality dataset adapted from real service interactions and refined for structured strategy use. To address the lack of high-quality training data, we also develop a role-playing approach that creates RoleCS, a synthetic dataset of strategy-rich conversations by assigning LLMs distinct roles aligned with the CSC framework. Fine-tuning on RoleCS significantly boosts LLMs’ ability to generate strategy-aligned and effective responses on CSConv.
@@ 2 Related Work
The goal of CSConv, which is to generate conversations based on real spoken customer service interactions, differs significantly from previous task-oriented dialogue datasets. Many earlier synthetic datasets have been created using the Wizard of Oz (WOZ) framework (kelley-1984-an), where one person acts as the system and another as the user. wen-etal-2017-network introduce a crowdsourced version of the WOZ framework to collect domain-specific dialogue data more efficiently. Other task-oriented dialogue datasets include Frames (asri-etal-2017-frames), MultiWOZ (budzianowski-etal-2018-multiwoz), MultiDoGO (peskov-etal-2019-multi), EMPATHETICDIALOGUES (rashkin-etal-2019-towards), Taskmaster-1 (byrne-etal-2019-taskmaster), ESConv (liu-etal-2021-towards), and NATCS (gung-etal-2023-natcs), each of which explores different domains and annotation schemes for developing conversational agents.
@@ 2.2 Role-Playing in Conversation Generation
Recent advances in LLMs have enabled the use of role-playing agents for conversation generation. For example, bae-etal-2022-building generate dialogues aligned with specific roles, while yang-etal-2024-crafting simulate characters and settings to produce coherent interactions. wu-etal-2024-from create interactive narrative dramas, and ye-etal-2025-sweetiechat focus on emotional support conversations. Building on the ye-etal-2025-sweetiechat, we design role-playing agents for customer support with clearer role definitions and responsibilities to more closely reflect real-world service interactions.
@@ 3.1 CSC Framework
Building on the three core stages of supportive communication from hill-2019-help and the COPC practical guidelines, we work with domain experts to refine and extend this structure for customer support. The resulting CSC framework defines five stages: Connecting (greeting and rapport building), Identifying (understanding the customer’s issue and emotional state), Exploring (discussing and evaluating potential solutions), Resolving (delivering and confirming resolution), and Maintaining (closing the interaction while preserving the customer relationship). Figure 2 illustrates the flow across these stages.
@@ 3.2 Dataset Construction
Appendix B provides details on filtering rules, prompt templates, and annotation guidelines. For rewriting, we use DeepSeek-R1 (guo-etal-2025-deepseek-r1) instead of GPT-4o (openai-2024-gpt4o), as the latter tends to produce shorter, less emotionally rich dialogues. See Appendix B.5 for a comparison.
@@ 3.4 CSC Task Definition
Following the task of emotional support conversation (liu-etal-2021-towards; ye-etal-2025-sweetiechat), we define the CSC task as generating the supporter’s response. Specifically, at turn $k$, where $P_{k}=S$, the model receives the conversation history $X_{k}=\{\left(P_{i},T_{i},U_{i}\right)\}_{i=1}^{k-1}$ as input. The CSC task is then divided into two subtasks: [^5] We further investigate how strategy prediction contributes to improving response generation in Section 6.3.
@@ 4 Synthetic Conversation Generation with Role-Playing Agents
Fine-tuning LLMs has shown effective for improving performance in task-specific dialogue applications. However, building high-quality, multi-turn customer support datasets remains challenging due to the need for domain expertise, contextual coherence, and dialogue diversity. This limits both the scalability and scenario coverage. While recent studies use LLMs for dataset augmentation via rewriting or imitation, they often yield limited variation in dialogue flow and support strategy. To address this, inspired by ye-etal-2025-sweetiechat, we adopt a multi-role role-playing framework to generate synthetic customer support dialogues that are diverse, coherent, and representative of real-world scenarios.
@@ 4.1 Role Playing Conversation Generation
As shown in Figure 4, our role-playing framework involves five roles: Planner, Supporter Assistant, Supporter, Customer Assistant, and Customer. Each role serves a distinct purpose. The Planner defines the dialogue scenario and sets the Customer’s communication goal. The main interaction occurs between the Supporter and the Customer, simulating a real customer service exchange. Meanwhile, their respective Assistants, the Supporter Assistant and the Customer Assistant, offer strategic guidance to help them fulfill their roles more effectively. We use deepseek-r1 (guo-etal-2025-deepseek-r1) to simulate all roles, with detailed prompts available in Appendix E. We describe each role below.
@@
Diversity in conversations relies heavily on the richness of customer personas (wang-etal-2025-opencharacter). To support this, we construct a rich character profile pool $O$ that captures a wide spectrum of customer backgrounds, behaviors, and communication styles, thereby enhancing the realism and variability of the simulated interactions.
@@ 5.1 Experimental Settings
We evaluate several widely used LLMs on the CSC task, including GPT-4o (openai-2024-gpt4o), DeepSeek-R1 (guo-etal-2025-deepseek-r1), DeepSeek-V3 (liu-etal-2024-deepseek-v3), Qwen-2.5-7B/72B-Instruct (yang-etal-2024-qwen2.5), and LLaMA-3.1-8B/70B-Instruct (grattafiori-etal-2024-llama3). To evaluate the impact of the proposed RoleCS, we fine-tune the Qwen and LLaMA models using 137,406 fine-tuning instances extracted from RoleCS, formatted consistently with our single-prompt in test-time. Fine-tuning and inference configurations are provided in Appendix J.
@@
Following prior work (liu-etal-2021-towards; ye-etal-2025-sweetiechat), we adopt a diverse set of metrics to assess the response quality. These include BLEU-n (B-n(papineni-etal-2002-bleu) and ROUGE-L (R-L(lin-2004-rouge) for measuring lexical overlap, BERTScore (BS(zhang-etal-2020-bertscore) and BLEURT (BR(sellam-etal-2020-bleurt) for semantic similarity. In addition, to evaluate alignment with the intended support strategy, which is a core aspect of the CSC task, we report strategy prediction accuracy (ACC).
@@ 5.2 Experimental Results
  • Evaluation with generated context shows similar performance trends as with reference context, but with lower absolute scores. This drop reflects the difficulty of maintaining consistency and quality in multi-turn conversations when relying on model-generated dialogue history, echoing findings in prior work such as ye-etal-2025-sweetiechat.
@@ 6.4 Evaluation with LLMs and Human as Judges
For human evaluation, we randomly select 100 conversations and have them independently rated by three professional annotators on a 1-5 Likert scale (joshi-2015-likert) across the same six dimensions. Table 6 shows the average overall scores, which align with trends from the LLM-based evaluations. Table 7 reports Fleiss’ Kappa (fleiss-1971-kappa) scores among annotators and between GPT-4o-Judge and the annotators. The results show strong inter‑rater agreement and confirm that GPT-4o-Judge aligns well with human judgment.
@@ Appendix H Diversity Analysis
We examine the diversity of the CSConv and RoleCS datasets from two perspectives: lexical diversity and semantic diversity. To assess lexical diversity, we adopt Distinct-n (li-etal-2016-diversity) metric. For a fair comparison, we randomly select 1,000 dialogues from each dataset, removing speaker and strategy information, and utilizing Qwen2.5-7B tokenizer for tokenization. Regarding semantic diversity, we calculate cosine similarity between pairs of distinct dialogues using TF-IDF features (salton-1975-vector). Table 12 and Figure 28 present the results, respectively.
@@ Appendix I Prompt for Predicting Strategy and Generating Response
For fine-tuning, we use the LLaMA-Factory framework (zheng-etal-2024-llamafactory), applying LoRA (hu-etal-2022-lora) with a rank of 8 and a scaling factor of 16. All fine-tunings are conducted on 4 NVIDIA A100 80GB GPUs, with a batch size of 4 and gradient accumulation over 2 steps. The learning rate is set to 3e-5, and fine-tuning is run for 3 epochs to mitigate the risk of overfitting.